Project

General

Profile

2018-01-23-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

23rd January 2018; 15:30 GMT
RAL, R1 2nd Floor East, MICE Meeting room

Phone details

Present: KL, CW, CR, DR, AK, ABr, MP, PG

  1. Introduction: KL
    • Meeting will cover, as last time, decommissioning, work in the Hall, field survey and analysis.
  2. Notes of previous meeting and status of actions: All
    • Notes accepted. Status of actions:
      • CW: Reinstate the QD/QP systems.
        • Done. System re-instated. Need now to organise a test. Action CW.
      • CW: Prepare plan to carry out a magnetic-field survey for discussion.
        • Done. CERN field-mapping colleagues contacted. Do not (yet) have agreement with supervisor. Need now to re-contact the supervisor (KL). If this is OK, then need to decide to execute the field survey in leue of a pressing need to take data. The mechanical survey in ISIS sync. and DSA is done. Now need to measure Q7--9, TOF, KL and s/c magnets. Need to purchase equipment so can start moving detectors etc. once survey is done. Email to MICE has been sent to solicit input in advance of movement of equipment. Plan to advance movement of equipment as the necessary operations are time-consuming.
    • CR: Maintain the 'requirements' document as necessary.
      • Ongoing. Additional suggestion: measurement of absorber; e.g. weight measurement. Weight measurement is problematic because water ingress into the LiH is on the edge. Agree: visual inspection when ABr is at RAL to bag up the LiH and to weigh on receipt at FNAL.
    • KL/CW: Make a step in the preparation of the h/w papers.
      • Ongoing.: Some sections of LH2 have been completed. Now pushing to complete the first full draft. H/w system paper has been structured. A skeleton will be circulated soon. In tehfuture take this item in PM's report.
  3. Memo on measurements required before experiment is decommissioned: CR
    • No update; points noted above.
  4. Hall status and stability of systems: CW
    • Ambient temperature in the Hall was rising with concomitent effect on the FC pressure. The ambient temperature in the Hall has been corrected.
  5. Planning for decommissioning and making safe to be complete by 31Mar18: CW
    • KL/CW: Issue first version of the decommissioning plan this week.
    • Need to agree with Andy Nichols how to make the sign-off, draft exists.
  6. Software and computing status: DR
    • See notes at link above. Points noted:
      • Tracker-efficiency review will be reconvended once reprocessing of data has been analysed by C.Hunt.
      • Geometry: the LH2 geometry was not correct, this is now corrected. There was also an error in the LiH absorber geometry which has also been fixed.
      • Processing for emittance paper has been completed.
  7. Data validation ahead of start of Cycle 2017/04: CR
    • V.Pec and S.Wylbur are checking through the ToF data and the tracker group is ready to sign-off with the tracker data. CR has sampled the data using standard physics-check plots. Only issue seems to be that of pile-up in TOF0. CR will check once more and then pass issue to TOF0. Presently not believed to be a problem.
    • Also looking at doing a comprehensive analysis on 2017/02 and 2017/03. This will be a more detailed check.
    • CR amd DR will report to KL. Negotiate time and place offline. Presumption is we're going ahead with the preparations for the magnetic-field survey.
  8. Paper status and plans: CR
    • Emittance measurement paper:
      • Wisepeople comments generally minor. Systematic effects seem at present to be of the order of the statistical error, but, more MC data is required and further analysis is underway.
    • Emittance evolution rapid communication:
      • In progress. Slides presented at analysis w/s.
    • Field-off scattering:
      • Now has diffuser cut, improved data selection. Correction to selection and systematic errors.
    • Field-on scattering:
      • Seems to be stalled; need to add support for authors.
    • System-performance paper:
      • Seems to be stalled; need to add support for authors.
    • All the other data (including the wedge) analysis is underway.
  9. AOB
    • CERN Courier: need science and politix.

Summary of actions:

  • CW: Organise test of QD/QP systems for SSs.
  • KL: Contact supervisor of CERN field-mapping team to seek to secure permission for them to carry out the MICE magnetic-field survey.
  • KL/CW: Issue first version of the decommissioning plan this week.

Back to MEMO:

Back

Back to MIPO:

Back

Documents